Free OAT Exam Braindumps
Optometry Admission Questions 81–88 (page: 12)

Updated On: 20-Jun-2026

For the following equation: 2x2 − 10x + 12 = 0, find the roots.

  1. x = 1, x = 4
  2. x = 2, x = 6
  3. x = 0, x = 1
  4. x = 2, x = 3
  5. None of the above

Answer(s): D



The difference between one half of a number and one eighth of a number is 15. What is this number?

  1. 32
  2. 40
  3. 48
  4. 56
  5. 64

Answer(s): B

Explanation:

Start with 32. If you half it, that is 16. One eighth of 32 is 4. Thus the difference between these numbers is 12.
Proceed to 40. Half of 40 = 20. One eighth of 40 is 5. Thus, the difference is 20 − 5 = 15, which is the answer.



A square has an area of 36. If you doubled the length of the sides of the square, what will its new perimeter be?

  1. 24
  2. 36
  3. 48
  4. 52
  5. 72

Answer(s): C

Explanation:

If a square has an area of 36, it’s sides must be 6, as s2 = area of a square. Doubling the side of the square gives 12. Add all 4 sides (12 + 12 + 12 + 12) and you have a perimeter of 48.



Sin(30+50) = ?

  1. cos30sin50 − sin30cos50
  2. cos80
  3. sin30cos50 + cos30sin50
  4. tan80 − cos30
  5. 2sin30cos50

Answer(s): C

Explanation:

Sin(x + y) = sinxcosy + cosxsiny



With a standard deck, what are the odds of drawing a 8 of hearts followed by a 2 of diamonds, without replacement?

  1. 1/2652
  2. 1/2704
  3. 1/50
  4. 1/3042
  5. 1/1064

Answer(s): A

Explanation:

The chance of drawing a 8 of hearts in a standard deck (52 cards) is simply 1/52. There is only one 8 of hearts in a deck. Without replacement of the 8 of hearts, the chance of drawing a 2 of diamonds is 1/51 (again, 51 because we did not put the 8 of hearts back in the deck, and there is only one 2 of diamonds in a deck). These 2 events are then multiplied because the questions asks for the chance of BOTH happening, not EITHER happening.
Thus: 1/52 × 1/51 = 1/2652



134 is what percent of 335?

  1. 25
  2. 35
  3. 40
  4. 45
  5. 55

Answer(s): C

Explanation:

For this (and other similar problems), it is basically asking, what percent of 335 will give 134. From the answer choices, you can already eliminate anything over 50% as 134 is clearly less than half of 335, and you can tell 134 will definitely fit into 335 less than 4 or so times, thus eliminating 25%. This leaves you with 3 answer choices.
Begin the process of elimination by multiplying the answer choice percentages by the given 335 and pick the one that gives the provided benchmark (134).
Answer choice (35) will give: 0.35 × 335 = 117.25. Answer choice (40) will give: 0.40 × 335 = 134



Juliet gets 78, 84, 98, 94 on her first 4 genetics exam. What grade must she receive on her final and fifth exam, in order to attain a 'A' in the class?

  1. 92
  2. 93
  3. 94
  4. 95
  5. 96

Answer(s): E

Explanation:

Make this an algebra type problem as you find the averages of the exam scores to total 450. (Because 450/500 is 90% = A)
78 + 84 + 98 + 94 + x = 450
Solving for x: 354 + x = 450 x = 96



A store has a series of sales. In August, all items are reduced to 50%. In September, all items are reduced an additional 30%. If an item was $140 back in July, how much would the item be in September?

  1. $49
  2. $70
  3. $35
  4. $30
  5. $77

Answer(s): A

Explanation:

Starting out with $140, taking 50% off (or simply doing 140 × 0.5) gives us $70. Taking an ADDITIONAL 30% off $70, we do 70 × 0.3=21. 70 − 21 = $49.



Viewing page 12 of 36
Viewing questions 89 - 96 out of 274 questions


Post your Comments and Discuss Test Prep OAT exam prep with other Community members:

AI Tutor AI Tutor 👋 I’m here to help!